# EpidemIQs: Prompt-to-Paper LLM Agents for Epidemic Modeling and Analysis

### v1

@@ Abstract
Large Language Models (LLMs) offer new opportunities to automate complex interdisciplinary research domains. Epidemic modeling, characterized by its complexity and reliance on network science, dynamical systems, epidemiology, and stochastic simulations, represents a prime candidate for leveraging LLM-driven automation. We introduce EpidemIQs, a novel multi-agent LLM framework that integrates user inputs and autonomously conducts literature review, analytical derivation, network modeling, mechanistic modeling, stochastic simulations, data visualization and analysis, and finally documentation of findings in a structured manuscript. We introduced two types of agents: a scientist agent for planning, coordination, reflection, and generation of final results, and a task-expert agent to focus exclusively on one specific duty serving as a tool to the scientist agent. The framework consistently generated complete reports in scientific article format. Specifically, using GPT 4.1 and GPT 4.1 mini as backbone LLMs for scientist and task-expert agents, respectively, the autonomous process completed with average total token usage 870K at a cost of about \$1.57 per study, achieving a 100% completion success rate through our experiments. We evaluate EpidemIQs across different epidemic scenarios, measuring computational cost, completion success rate, and AI and human expert reviews of generated reports. We compare EpidemIQs to the single-agent LLM, which has the same system prompts and tools, iteratively planning, invoking tools, and revising outputs until task completion. The comparison shows consistently higher performance of the proposed framework across five different scenarios. EpidemIQs represents a step forward in accelerating scientific research by significantly reducing costs and turnaround time of discovery processes, and enhancing accessibility to advanced modeling tools.
@@ I Introduction
Scientists’ Efforts to automate research date back to at least the early 1970s [1, 2, 3], aiming to overcome limits on the number of ideas scientists can pursue, constraints that force them to prioritize ideas with higher predicted impact. For example, the Automated Mathematician was introduced to carry out simple mathematics research guided by large sets of heuristic rules [4, 5], and DENDRAL was proposed to assist chemists with data interpretation problems [6].
More recently, advances in autoregressive large language models (LLMs) [7, 8, 9] have demonstrated strong capabilities in question answering, problem-solving, and coding, with promising applications across science and engineering. However, LLMs still face limitations when applied to real-world tasks [10]. These constraints led to the development of LLM agents, which extend LLMs with abilities such as external tool use [11, 12, 13], chain-of-thought prompting [14], and iterative self-improvement [15, 16]. LLM agents have since been applied to diverse domains, including software engineering [17, 18], cybersecurity [19], medical diagnosis [20], chemistry [21, 22, 23], material science [24, 25, 26], computational biology [27], algorithm design [28], and healthcare [29]. Multi-agent systems have also emerged to simulate human-like research environments across disciplines. Examples include Agent Laboratory [30], Virtual Lab [31], ChemCrow [32], ResearchAgent [33], and The AI Scientist [34]. For broader surveys of LLM agent applications, see [35, 36, 37, 38].
@@
Addressing these models requires an interdisciplinary expertise: (i) stochastic-process theory to analyze the master equations and their approximations; (ii) network science to characterize and parametrize contact structures; (iii) epidemiology to constrain pathogen-specific parameters; and (iv) computational epidemiology to conduct simulations and scenario analysis. Having a research group with this level of diversity is not only challenging to access but also poses coordination challenges. In this study, we introduce EpidemIQs, a novel multi-LLM agents capable of autonomously addressing research in network-based epidemic modeling for the first time.
EpidemIQs emulates an interdisciplinary research environment with five collaborative teams of task-expert and scientist agents, capable of performing end-to-end autonomous research: given a user query, if identified as a network-based problem, it initiates the process by discovering relevant insights from diverse online sources (e.g., web-based content and published papers) to conduct a literature review and address analytical aspects through mathematical reasoning. It then formulates the problem as a network-based model, tests it through rigorous stochastic simulations, analyzes the resulting multimodal data (e.g., visual and numerical), and finally composes a scientifically formatted manuscript that reports the entire procedure, from problem formulation to simulation results, discussion, and final conclusion. The framework can also operate in copilot mode, which enables human cooperation throughout the process.
We evaluate our proposed framework across five epidemic scenarios articulated through questions, three of which were unknown to the framework. We tested EpidemIQs against the multi-turn single-agent baseline, comparing performance by (i) discussing the results of each team across all scenarios, (ii) AI and human expert evaluations of generated papers, and (iii) computational statistics. The proposed framework consistently showed impressive performance, achieving a 100% completion success rate, an average human review score of 7.98 out of 10, a low cost of only \$1.57, and a total processing time of 1,190 seconds. We can summarize the main contributions of our work as follows:
@@
  • Enabled fully autonomous mode to minimize human intervention, requiring only an initial query, while also supporting a copilot mode for collaborative operation.
  • Ensured high interpretability by having each scientist provide reasoning for its actions and generated outcomes, and recording results and scripts to facilitate validation of procedures and outcomes.
The remainder of this paper is organized as follows: Section II describes the methodology. Section III details the design and implementation of the system. Section IV outlines the experimental setup and evaluation cases. Section V presents the results and discusses the study’s limitations and shortcomings. Finally, Section VII concludes the paper and proposes directions for future work.
@@ II Methodology
LLMs demonstrate advanced capabilities relevant to scientific tasks, including coding and answering technical questions. However, their performance in tackling complex epidemiology scenarios remains limited. To address these challenges, we propose a multi-agent framework that emulates a scientific epidemic research laboratory environment, designed to enhance the practical capabilities of LLMs in complex scientific tasks. The EpidemIQs framework is organized into four core functional layers that collectively support autonomous research workflows[38]: (1) The multi-agent orchestration layer governs the coordination, communication, and cooperation among specialized agents.(2) The backbone LLM operates as the central reasoning and decision-making unit . (3) The perception layer collects, integrates, and processes heterogeneous data sources, including scientific literature, experimental results, contextual information, and prior knowledge bases. (4) The action layer executes tasks derived from the LLM’s output, encompassing activities such as code generation, simulation, modeling, etc.
Figure 1: Multimodal approach for the multi-agent framework for epidemic autonomous research. This framework enhances the power of multi-LLM agents by integrating multimodal data from various sources, such as literature, online web data, theoretical models, and simulation and experiments.
Looking at Figure 1, the architectural design illustrates the structured interaction among the four functional layers that enable autonomous epidemic research. The orchestration layer governs inter-agent communication, task allocation, and coordination protocols, ensuring efficient collaboration among specialized agents. Within the perception layer, agents process multimodal data streams through standardized interfaces that convert heterogeneous sources, including scientific literature, experimental data, contextual metadata, and structured knowledge bases, into unified semantic representations suitable for downstream reasoning and analysis. The backbone LLM operates as the central cognitive engine, analyzing perceived inputs, generating hypotheses, formulating research plans, and executing decisions based on accumulated evidence and domain constraints. The agents’ outputs inform the action layer, where the system autonomously selects and calls functions to deploy tools, such as stochastic simulation engines, Retrieval-Augmented Generation (RAG), API calls, and executes it in the correct order, so that high‑level research goals are broken down into concrete, reproducible steps such as loading data, running simulations, and storing results.
@@ III Multi-LLM Agent Architecture: EpidemIQs Framework
The orchestration of EpidemIQs is illustrated in Figure 2. This framework is structured to emulate a scientific laboratory, comprising five phases: Discovery, Modeling, Simulation, Analysis, and Report Writing. These phases are executed by distinct collaborative multi-agent system teams, and the results of their work are integrated to produce the final report. In this section, we first describe the design of two types of agents, followed by the architecture design, output structure, memory, and implementation details of EpidemIQs.
@@
III-A Two Agent Types: Scientists and Experts
Two agent types are inspired by the generative agent concept [42] and further advanced by frameworks such as CAMEL [43], AutoGen [44], and Metagpt [45], and are designed to enable autonomous scientific reasoning and investigation. The central coordinating component, referred to as the scientist agent, performs high-level orchestration through iterative planning, reflection, and execution refinement. Upon receiving an input query, the scientist employs a dedicated plan module, which parses and decomposes the query into a plan comprising sub-tasks and their logical dependencies. Both reflect and ReAct modules are constrained to generate output in predefined formats. While structured outputs, e.g., JSON, enhance reproducibility and reliability of performance, they have recently been shown to compromise the reasoning capability of LLMs [46]. Therefore, we explicitly devote a separate plan module as an agent to allow it to freely generate the plan, while both the subsequent ReAct and reflect modules are restricted to generating predefined structured outputs.
Once the plan is formulated, the scientist agent iteratively executes the ReAct-Reflect loop until either the reflect agent decides no further revision is needed, or it reaches the maximum iteration limit, $r_{max}$. The ReAct Module executes reasoning, and call tools conditioned on the current plan segment, and refine its response based on the action outcomes, while the reflect module prompted to perform logical reasoning, evaluates intermediate outputs for errors, logical inconsistencies, or incomplete results, producing structured JSON feedback to refine action selection and output validity across iterations, enhancing the agent performance through linguistic feedback without need to update the weights of models.
In parallel, the scientist agent can coordinate multiple independent task-expert agents as tools, each designed as a specialized ReAct agent for a specific task or data source. These expert agents serve as domain specialists responsible for well-defined atomic tasks such as literature retrieval, online search, mathematical derivation, or data extraction. This functional decomposition ensures that each expert executes its task with minimal context switching, while the scientist maintains global task coherence and quality control.
III-B Multi-Agent Orchestration
The EpidemIQs framework is comprised of 5 phases. It initiates upon receiving a user query, entering the Discovery phase. In this phase, the DiscoveryScientist agent performs planning and reasoning to collect data relevant to the query following the Multi-Hop Question Answering (MHQA) paradigm [47]. Specifically, the data collection process is conducted in three iterative steps, with each step focusing on a distinct information source. This multi-step procedure enables the system to progressively refine its understanding, reduce errors, and enhance the overall accuracy of the gathered information[48]. The components of the Discovery team are as follows:
  • DiscoveryScientist coordinates the expert agents to collect relevant information, process the retrieved data, and synthesize the final output following a predefined structured format. This output encompasses a comprehensive description of the scenario, task specifications, and contextual information related to the disease and the affected population.
  • OnlineRetrieverExpert is responsible for acquiring relevant information from online sources. It generates and refines search queries in multiple stages, leveraging external APIs (e.g., Tavily API) to retrieve data from the web. The agent subsequently processes and integrates the retrieved information to generate a coherent response.
  • LiteratureRetrieverExpert conducts a multi-step search process using the Semantic Scholar API, based on the received query. For each retrieved query, the agent generates concise summaries and aggregates the most pertinent findings. The agent is prompted to generate the final summarized response based on multiple criteria, including relevance and accuracy.
  • MathExpert provides mathematical reasoning and analytical solutions for queries requiring quantitative analysis. The agent is capable of both generating and executing Python code, a capability demonstrated to enhance solution accuracy and mitigate hallucination risks ( [49] showed that it can increase up to 39.8% problem-solving capability over HumanEvals [50] benchmark). The backbone LLM of this agent should be an advanced language model with high reasoning capacity, able to perform chain-of-thoughts( such as OpenAI GPT o3-Jan-2025 or o3-mini-Jan-2025 models [51])
Following the completion of the Discovery phase, the Modeling phase is initiated to systematically construct the foundational components required for simulating epidemic dynamics on complex networks. This phase is orchestrated by a team of three specialized scientist agents, each responsible for a distinct modeling dimension: network topology, mechanistic model formulation, and parameterization. The functional responsibilities of these agents are delineated as follows:
  • NetworkScientist is tasked with generating a representative contact network that encapsulates the interaction patterns within the target population. The constructed network serves as a structural substrate upon which the epidemic process unfolds. Depending on the characteristics of the population and the epidemiological scenario, the network may adopt various configurations, including but not limited to regular networks or Erdős-Rényi random graphs [52], Barabási-Albert scale-free networks [53], and multilayer structures [54]. The NetworkScientist employs algorithmic procedures implemented in Python, utilizing scientific computing libraries such as NetworkX and SciPy, to model the network topology. The finalized network structure is preserved in the workflow repository for integration into the simulation pipeline.
@@ III-B Multi-Agent Orchestration
  • ModelerScientist is responsible for defining the mechanistic framework of disease progression. It selects and configures an appropriate compartmental model based on the disease characteristics and insights derived during the Discovery phase.
  • ParameterScientist is invoked after the design of both network topology and compartmental model, to calculate and assign quantitative values to model parameters and initialize network states. It performs calculations through the Python coding tool.
Upon completion of the modeling phase, the framework transitions into the Simulation, wherein predictive experiments are conducted to analyze the modeled epidemic dynamics. This phase is executed by the SimulationScientist, which performs multiple stochastic simulations to capture the probabilistic evolution of contagion spread over the constructed contact network, and has delegated VisionExpert to verify successful outcomes of simulations. The VisionExpert has vision-language capabilities and is responsible for analyzing the visual outputs generated during the simulation phase. It interprets time-series plots or state evolution graphs, enabling interpretability and validation of dynamic behavior that may not be easily deduced from raw numerical outputs alone.
To conduct these simulations, the agent utilizes the Fast Generalized Epidemic Modeling Framework (FastGEMF) [55], a high-performance simulation engine capable of modeling arbitrary multi-compartment epidemic processes over multilayer network structures with computational efficiency. The SimulationScientist is pre-trained through one-shot learning examples to interface with FastGEMF. Throughout the simulation process, the agent records the generated time-series data using Python libraries such as Pandas and stores both the numerical outputs and corresponding state-evolution visualizations in the project repository.
Following the simulation stage, the framework advances to the Analysis, wherein a coordinated team of specialized agents derives actionable insights from the simulation outputs. This multi-agent analysis team is responsible for processing both numerical results and visual artifacts to extract context-specific metrics and interpret spread dynamics. The roles and responsibilities of the constituent agents are defined as follows:
  • DataScientist coordinates the specialized expert agents and integrates their outputs into a coherent interpretation of the simulation results. It identifies and computes key epidemiological metrics, such as peak infection time and size, epidemic duration, outbreak probability, etc., tailored to the nature of the modeled disease and the research query. The final output consists of numerical measures and qualitative analysis.
  • VisionExpert analyzes visualizations to extract metrics and interpret epidemic behavior (e.g., uncertainty bands around trajectories), that may not be easily interpretable from raw numerical data.
  • DataExpert is tasked with the interpretation of structured numerical data. This agent accesses simulation results stored in the system repository and applies analytical techniques using Python libraries such as Pandas, NumPy, and SciPy. Operating within a multi-hop paradigm, the DataExpert initially identifies the structure and content of the available data, then dynamically selects appropriate analysis strategies to compute requested metrics. This flexibility ensures the agent’s ability to generalize across varying data types and experimental settings.
The final phase of the framework is Automated Report Generation, wherein two coordinated expert agents synthesize the research findings into a structured academic manuscript. This process starts with the ReporterExpert, which is equipped with tools to access the outputs of the entire workflow, including simulation results, code scripts, and literature summaries. Guided by a predefined LaTeX-based manuscript template, the ReporterExpert sequentially drafts all major sections of the paper, such as the Title, Abstract, Introduction, Methods, Results, and Discussion. Following the generation of each section, the LatexCraftExpert is tasked to perform a syntactic and structural validation of the LaTeX code, resolving potential compilation errors and ensuring stylistic consistency with publication guidelines. Upon completion of the main draft, the LiteratureReviewExpert conducts a contextual relevance check, retrieving and integrating recent scholarly works aligned with the study’s findings. This agent augments the manuscript with a Background and Related Work section to strengthen the research academic positioning. Once the initial draft is completed, an iterative review loop is repeated by the ReporterExpert. It critiques each section to detect inconsistencies, redundancies, or logical gaps according to a user-defined number of revision cycles, $n_{max}$. Finally, it is compiled as a PDF file.

### v2

@@ Abstract
Large Language Models (LLMs) offer new opportunities to accelerate complex interdisciplinary research domains through AI assistant tools. Epidemic modeling, characterized by its complexity and reliance on network science, dynamical systems, epidemiology, and stochastic simulations, represents a prime candidate for leveraging LLM-driven automation. We introduce EpidemIQs, a novel multi-agent LLM framework that integrates user inputs and autonomously conducts literature review, analytical derivation, network modeling, mechanistic modeling, stochastic simulations, data visualization and analysis, and finally documentation of findings in a structured manuscript, through five predefined research phases. We introduce two types of agents: a scientist agent for planning, coordination, reflection, and generation of final results, and a task-expert agent to focus exclusively on one specific duty serving as a tool to the scientist agent. The framework consistently generated complete reports in scientific article format. Specifically, using GPT 4.1 and GPT 4.1 Mini as backbone LLMs for scientist and task-expert agents, respectively, the autonomous process completes with average total token usage 870K at a cost of about \$1.57 per study, successfully executing all phases and final report. We evaluate EpidemIQs across several different epidemic scenarios, measuring computational cost, workflow reliability, task success rate, and LLM-as-Judge and human expert reviews to estimate the overall quality and technical correctness of the generated results. Through our experiments, the framework consistently addresses evaluation scenarios with an average task success rate of 79%. We compare EpidemIQs to an iterative single-agent LLM, benefiting from the same system prompts and tools, iteratively planning, invoking tools, and revising outputs until task completion. The qualitative and quantitative comparisons suggest a consistently higher performance of the proposed framework across the tested scenarios. We discuss the benefits that EpidemIQs can offer, its limitations, and ethical concerns of such AI research assistant tools.
@@ I Introduction
Scientists’ efforts to automate research date back to at least the early 1970s[1, 2, 3], aiming to overcome limits on the number of ideas scientists can pursue, constraints that force them to prioritize ideas with higher predicted impact [4, 5]. For example, the Automated Mathematician was introduced to carry out simple mathematics research guided by large sets of heuristic rules [4, 5], and DENDRAL was proposed to assist chemists with data interpretation problems [6].
More recently, advances in autoregressive large language models (LLMs) [7, 8, 9] have demonstrated strong capabilities in question answering, problem-solving, and coding, with promising applications across science and engineering. However, LLMs still face limitations when applied to real-world tasks [10]. These constraints led to the development of LLM agents, which extend LLMs with abilities such as external tool use [11, 12], chain-of-thought prompting [13], and iterative self-improvement [14]. LLM agents have since been applied to diverse domains, including software engineering [15, 16], cybersecurity [17], medical diagnosis [18], chemistry [19, 20], material science [21, 22, 23], computational biology [24], algorithm design [25], and healthcare [26]. Multi-agent systems have also emerged to simulate human-like research environments across disciplines. Examples include Agent Laboratory [27], Virtual Lab [28], ChemCrow [29], ResearchAgent [30], and The AI Scientist [31]. For broader surveys of LLM agent applications, see [32].
@@
Addressing these models requires an interdisciplinary expertise: (i) stochastic-process theory to analyze the master equations and their approximations; (ii) network science to characterize and parametrize contact structures; (iii) epidemiology to constrain pathogen-specific parameters; and (iv) computational epidemiology to conduct simulations and scenario analysis. Having a research group with this level of diversity is not only challenging to access but also poses coordination challenges. In this study, we introduce EpidemIQs, a novel multi-LLM agents system that can execute an end-to-end workflow to address network-based epidemic modeling with minimal human intervention.
EpidemIQs emulates an interdisciplinary research environment with five collaborative teams of task-expert and scientist agents, capable of performing end-to-end workflow autonomously: given a user query, if identified as a network-based problem, it initiates the process by discovering relevant insights from diverse online sources (e.g., web-based content and published papers) to conduct a literature review and address analytical aspects through mathematical reasoning. It then formulates the problem as a network-based model, tests it through rigorous stochastic simulations, analyzes the resulting multimodal data (e.g., visual and numerical), and finally composes a scientifically formatted manuscript that reports the entire procedure, from problem formulation to simulation results, discussion, and final conclusion. The framework can also operate in copilot mode, which enables human cooperation throughout the process.
To demonstrate the capability, we test our proposed framework across ten scenarios, focusing on five main aspects of epidemic scenarios articulated through questions, three of which were unknown to the framework. Other scenarios, including real-world outbreak data, opinion dynamics, and financial contagion, that share similar dynamics are presented in the Supplementary Material. We tested EpidemIQs against the iterative single-agent as a baseline for ablation of orchestration, comparing performance by (i) discussing the results of each team across all scenarios, (ii) AI and human expert evaluations of generated results and papers, (iii) workflow completion and task success rates, and (iv) computational statistics. The proposed framework consistently showed impressive performance, achieving a 100% workflow completion rate and 79±8.6% task success rate, an average human review score of 7.98±0.2 out of 10, a low cost of only \$1.57, and a total turnaround time of 1,190 seconds. We can summarize the main contributions of our work as follows:
@@
  • Enabled fully autonomous mode to minimize human intervention for workflow completion, requiring only an initial query, while also supporting a copilot mode for collaborative operation.
  • Ensured high interpretability by having each scientist agent provide reasoning for its actions and generated outcomes, and recording results and scripts to facilitate validation of procedures and outcomes.
The remainder of this paper is organized as follows: Section II describes the methodology. Section III details the design and implementation of the system. Section IV outlines the experimental setup, evaluation cases, and metrics. Section V presents the ablation study, results, and discusses the study’s limitations and shortcomings. Finally, Section VII concludes the paper and proposes directions for future work.
@@ II Methodology
LLMs demonstrate advanced capabilities relevant to scientific tasks, including coding and answering technical questions. However, their performance in tackling complex epidemiology scenarios remains limited. To address these challenges, we propose a multi-agent framework that emulates a scientific epidemic research laboratory environment to enhance the practical capabilities of LLMs for complex scientific tasks. The EpidemIQs framework is organized into four core functional layers that collectively support autonomous workflow[32]: (i) The multi-agent orchestration layer. (ii) The backbone LLM . (iii) The perception layer . (iv) The action layer executes tasks derived from the LLM’s output, encompassing activities such as code generation, simulation, modeling, etc.
Figure 1 illustrates how these layers interact to support an end-to-end workflow. The orchestration layer assigns tasks and synchronizes agent interactions so that different subtasks are handled coherently. The perception layer converts diverse inputs into unified semantic representations that can be reasoned over. Based on these inputs, the backbone LLM analyzes the evidence, forms hypotheses, and determines the sequence of actions required to address the research question. The action layer then autonomously calls appropriate tools, such as stochastic simulators, Retrieval-Augmented Generation (RAG), API calls, and executes them in the correct order.
Figure 1: Multimodal approach for the multi-agent framework for flexible epidemic research. This framework enhances the power of multi-LLM agents by integrating multimodal data from various sources, such as literature, online web data, theoretical models, and simulation and experiments.
Through this process, the system translates a high level research question into executable steps and produces results that directly address the posed objective.
@@ III Multi-LLM Agent Architecture: EpidemIQs Framework
The orchestration of EpidemIQs is structured to emulate a scientific laboratory, comprising five phases: Discovery, Modeling, Simulation, Analysis, and Report Writing. These phases are executed by distinct collaborative multi-agent system teams, and the results of their work are integrated to produce the final report. In this section, we first describe the design of two types of agents, followed by the architecture design, output structure, memory, and implementation details of EpidemIQs.
III-A Two Agent Types: Scientists and Experts
Two agent types are inspired by the generative agent concept [35] and further advanced by frameworks such as CAMEL [36], AutoGen [37], and Metagpt [38], and are designed to enable autonomous scientific reasoning and investigation. The central coordinating component, referred to as the scientist agent, performs high-level orchestration through iterative planning, reflection, and refinement of execution, as shown in Figure 2. Upon receiving an input query, the scientist agent employs a dedicated plan module, which parses and decomposes the query into a plan comprising sub-tasks and their logical dependencies. Both reflect and ReAct modules are constrained to generate output in predefined formats. While structured outputs, e.g., JSON, enhance reproducibility and reliability of performance, they have recently been shown to compromise the reasoning capability of LLMs [39]. Therefore, we explicitly devote a separate plan module as an agent to allow it to freely generate the plan, while both the subsequent ReAct and reflect modules are restricted to generating predefined structured outputs.
Once the plan is formulated, the scientist agent iteratively executes the ReAct-Reflect loop until either the reflect agent decides no further revision is needed, or it reaches the maximum iteration limit, $r_{max}$. The ReAct Module executes reasoning, and call tools conditioned on the current plan segment, and refine its response based on the action outcomes, while the reflect module prompted to perform logical reasoning, evaluates intermediate outputs for errors, logical inconsistencies, or incomplete results, producing structured JSON feedback to refine action selection and output validity across iterations, enhancing the agent performance through linguistic feedback without need to update the weights of models. In parallel, the scientist agent can coordinate multiple independent task-expert agents as tools, each designed as a specialized ReAct agent for a specific task or data source. These expert agents serve as domain specialists responsible for well-defined atomic tasks such as literature retrieval, online search, mathematical derivation, or data extraction. This functional decomposition ensures that each expert agent executes its task with minimal context switching, while the scientist agent maintains global task coherence and quality control.
III-B Multi-Agent Orchestration
@@
Figure 3 illustrates that EpidemIQ’s framework operates after receiving the user query, beginning with Discovery. Here, the Discovery Scientist utilizes a Multi-Hop Question Answering (MHQA) paradigm [40] across three iterative steps to progressively refine data collection accuracy [41]. The team comprises:
  • Discovery Scientist: Coordinates expert agents to synthesize a structured output defining the scenario, task specifications, and disease context.
  • Online Retriever Expert: Fetches and refines relevant data from web sources via external APIs (e.g., Tavily).
  • Literature Retriever Expert: Aggregates and summarizes pertinent academic findings via the Semantic Scholar API.
  • Math Expert: Provides quantitative analysis and Python-based execution to enhance accuracy [42], utilizing reasoning capable LLMs (e.g., OpenAI-o3-Mini).
Subsequently, the Modeling phase constructs the components for simulating epidemic dynamics through three specialized agents:
  • Network Scientist: Models the population contact network (e.g., random, scale-free, or multilayer [43]) using scientific libraries like NetworkX.
@@
  • Model Scientist: Defines the mechanistic disease progression framework by selecting appropriate compartmental models based on Discovery phase insights.
  • Parameter Scientist: Calculates and assigns quantitative model parameters and initial network states via Python code execution.
The framework then transitions to Simulation, executing stochastic experiments via the Fast Generalized Epidemic Modeling Framework (FastGEMF) [44]. The Simulation Scientist configures the engine for multi-compartment processes on multilayer networks, recording time-series data and state-evolution visualizations. The Vision Expert verifies simulation outcomes to ensure validity. In the Analysis phase, a multi-agent team derives actionable insights:
  • Data Scientist: Coordinates the team to compute epidemiological metrics (e.g., peak infection size, outbreak probability) and synthesize qualitative analysis.
  • Vision Expert: Interprets visual outputs (e.g., trajectory uncertainty bands) to validate dynamic behaviors not easily deduced from raw numbers.
  • Data Expert: Analyzes structured numerical data stored in the repository using Python libraries (Pandas, SciPy) via a multi-hop strategy.
Finally, Automated Report Generation synthesizes findings into an academic manuscript. The Reporter Expert drafts sections (e.g., Methods, Results) using a LaTeX template, while the Latex Craft Expert validates syntax and structure. The Literature Review Expert integrates related work for contextual positioning. The draft undergoes an iterative self-correction loop ($n_{max}$ cycles) before final PDF compilation. The details of the architecture are provided in Supplementary Section LABEL:S-sec:orch.